Solve for x and what is the common denominator? There should be a line under each top number. Thank you


1 + 3 = 2
4x 2x x+1

Answer by stanbon(75887)


1/(4x) + 3/(2x) = 2/(x+1)
---
lcd = 4x(x+1)
---
Multiply thru by lcd to get:
x+1 + 3*2(x+1) = 2*4x)
-----
x + 1 + 6x+6 = 8x
x = 7
----
Cheers,
Stan H.
=============